# Rigidity and nonexistence of CMC hypersurfaces in 5-manifolds

### v2

@@ 5. Proof of main results
5.1. Proof of Theorem 1.9
The conclusion follows from the argument in [chodosh-li-stryker] by constructing compactly supported test function. While, for completeness and independent interest (to prove Proposition 1.13), we provide the detailed construction of balanced test functions for weak stability. To do it, we need to decompose $M$ into convenient building-blocks (see [chodosh-li-stryker]).
Let $M^{4}$ be a complete, simply connected, Riemannian $4$-manifold such that the conclusions of Theorem 3.1 hold. Let $\{E_{l}\}$ be the nonparabolic end of $M$ adapted to $p\in M$ with length scale $L$.
We now decompose $M$ to deal with the nonparabolic end. Let $l_{0}\geqslant 1$. Then
@@
The third part aims to pack the parabolic ends attached to the nonparabolic end $E_{l}$.
@@
Note that for fixed $i\geq 1$, we have the decomposition
By Theorem 3.1, $P_{l}^{(\alpha)}$ is either bounded or $P_{l}^{(\alpha)}\backslash K$ is parabolic for all compact $K\subset M$ such that $P_{l}^{(\alpha)}\backslash K$ has smooth boundary. It is important to keep in mind that as $l\rightarrow\infty$, $n_{l}$ may not be bounded.
To prove the main theorem, we recall the following fact.
Proposition 5.1.
For any $l$ and $\alpha\in[1,n_{l}]$ and for any small $\epsilon>0$, there exists a positive compactly supported Lipschitz function $u_{l,\alpha}$ on $P_{l}^{(\alpha)}$ such that
$u_{l,\alpha}|_{\partial P_{l}^{(\alpha)}}=1\ \text{and}\ \ \int_{P_{l}^{(\alpha)}}|\nabla u_{l,\alpha}|^{2}<\epsilon.$
This follows from the definition of the parabolicity of ends. See [peterwangspectrum] or [chodosh-li-stryker, Proposition 3.4] for the detailed explanation.
Proof of Theorem 1.9.
Suppose that a two-sided weakly stable minimal hypersurface $M^{4}$ exists in $(X^{5},g)$ of Theorem 1.9, let $\{E_{l}\}$ be the nonparabolic end of $M$. We use the deocomposition of $M$ constructed above.
For each $l$, let $\rho_{l}$ be a smooth function on $\overline{M}_{l}$ such that $|\nabla\rho_{l}|\leqslant 2$,
For instance, we can take $\rho_{l}$ to be a smoothing of the distance function from $p$.
Since $M$ is weakly stable, outside a compact subset $\Omega$, $M$ is strongly stable. We can choose large $l_{0}$ such that $B_{l_{0}L}(p)$ contains $\Omega$. To prove the theorem, on the contrary we assume that $\omega(x)=|A_{M}|^{2}(x)+\operatorname{Ric}_{X}(\eta,\eta)$ is positive at some point $q$, then $\omega$ is positive in a neighborhood $\Omega_{q}$ of $q$ by continuity. We also assume that $\Omega_{q}$ is contained in $B_{l_{0}L}(p)$, this is achievable since $p$ and ł$\l_{0}$ are free to choose.
Let
$\int_{B_{l_{0}L}(p)}w\geq T$
where $T>0$ is fixed positive value.
Let $\phi_{1}:[l_{0}L,(l_{0}+2i)L]\to[-1,1]$ be the linear function defined as
$\phi_{1}(t)=1+\frac{1}{i}(l_{0}-\frac{t}{L}).$
Define a compactly supported Lipschitz function on $M$ as follows
$\varphi_{1}(x)=\begin{cases}1\ \ \ \ \ \ \ \ \ \ &x\in\overline{B}_{l_{0}L}(p),\\ \phi_{1}(\rho_{l}(x))&x\in\overline{M_{l}},\ \ \l_{0}\leq l<l_{0}+i,\\ \phi_{1}(lL)u_{l,\alpha}&x\in\overline{P}_{l}^{(\alpha)},\ \ l_{0}\leq l<l_{0}+i,\\ 0&\ \ \text{otherwise}.\end{cases}$
Define a Lipschitz function on $M$ as follows
$\varphi_{2}(x)=\begin{cases}0,\ \ \ \ \ \ \ \ \ \ &x\in\overline{B}_{l_{0}L}(p)\cup\left(\cup_{l=l_{0}}^{l_{0}+i-1}\overline{M_{l}}\right)\cup\left(\cup_{l=l_{0}}^{l_{0}+i-1}\overline{P}_{l}^{(\alpha)}\right),\\ \phi_{1}(\rho_{l}(x))&x\in\overline{M_{l}},\ \ \l_{0}+i\leq l<l_{0}+2i,\\ \phi_{1}(lL)u_{l,\alpha}&x\in\overline{P}_{l}^{(\alpha)},\ \ l_{0}+i\leq l<l_{0}+2i,\\ -u_{l,\alpha}&x\in\overline{P}_{l}^{(\alpha)},\ \ l_{0}+2i\leq l<l_{0}+3i+a,\\ -1&\text{otherwise}.\end{cases}$
Note that the two functions $\varphi_{1},\varphi_{2}$ have different supports on $M$.
Let $\phi_{2}:[(l_{0}+2i+a)L,(l_{0}+3i+a)L]\to[0,1]$ be the linear function defined as
$\phi_{2}(t)=-2+\frac{1}{i}(\frac{t}{L}-l_{0}-a).$
Define a Lipschitz function on $M$ as follows
$\varphi_{a}(x)=\begin{cases}0\ \ \ \ \ \ \ \ \ \ &x\in\overline{B}_{l_{0}L}(p)\cup\left(\cup_{l=l_{0}}^{l_{0}+2i+a-1}\overline{M_{l}}\right)\cup\left(\cup_{l=l_{0}}^{l_{0}+2i+a-1}\overline{P}_{l}^{(\alpha)}\right),\\ \phi_{2}(\rho_{l}(x))&x\in\overline{M_{l}},\ \ \l_{0}+2i+a\leq l<l_{0}+3i+a,\\ \phi_{2}(lL)u_{l,\alpha}&x\in\overline{P}_{l}^{(\alpha)},\ \ l_{0}+2i+a\leq l<l_{0}+3i+a,\\ 1&\ \ \text{otherwise}.\end{cases}$
Thus we can sum $\varphi_{2}$ and $\varphi_{a}$ to get a compactly supported Lispchtiz function. Let $\varphi_{s,a}=\varphi_{1}+s(\varphi_{2}+\varphi_{a})$ and we can easily see that $\varphi_{t}$ has compact support in $M$. We calculate
$\int_{M}\varphi_{1,a}$ $=\int_{M}\varphi_{1}+\int_{M}\left(\varphi_{2}+\varphi_{a}\right)$
$\leq\int_{M}\varphi_{1}-\sum_{l=l_{0}+2i}^{l_{0}+2i+a-1}\sum_{\alpha=1}^{n_{l}}\int_{P_{l}^{\alpha}}u_{l,\alpha}-\sum_{l=l_{0}+2i}^{l_{0}+2i+a-1}|M_{l}|.$
The first term of right hand side is fixed value. Thus we are able to choose large $a$ such that $\int_{M}\varphi_{1,a}<0$, due to the fact that $M$ has infinite volume. Denote such value by $a_{0}.$
Moreover, we know that $\int_{M}\varphi_{0,a_{0}}>0$. Since $\int_{M}\varphi_{s,a_{0}}$ is continuous in $s$, there exists $s_{0}\in(0,1)$ such that $\int_{M}\varphi_{s_{0}.a_{0}}=0$. Hence, it follows from the part regarding the interior ends of Lemma 4.8 and the direct computation that
$\int_{M}$ $|\nabla\varphi_{s_{0},a_{0}}|^{2}\leq\int_{M}|\nabla\varphi_{1}|^{2}+\int_{M}|\nabla\varphi_{2}|^{2}+\int_{M}|\nabla\varphi_{a_{0}}|^{2}+2\int_{M}\langle\nabla\varphi_{2},\nabla\varphi_{a_{0}}\rangle$
$=\sum_{l=l_{0}}^{l_{0}+2i-1}\int_{M_{l}}(\phi_{1}^{{}^{\prime}})^{2}|\nabla\rho_{l}|^{2}+\sum_{l=l_{0}+2i+a_{0}}^{l_{0}+3i+a_{0}-1}\int_{M_{l}}(\phi_{2}^{{}^{\prime}})^{2}|\nabla\rho_{l}|^{2}$
$+\sum_{l=l_{0}}^{l_{0}+2i-1}\sum_{\alpha=1}^{n_{l}}\phi_{1}^{2}(lL)\int_{P_{l}^{(\alpha)}}|\nabla u_{l,\alpha}|^{2}-2\sum_{l=l_{0}+2i+a_{0}}^{l_{0}+3i+a_{0}-1}\sum_{\alpha=1}^{n_{l}}\phi_{2}(lL)\int_{P_{l}^{(\alpha)}}|\nabla u_{l,\alpha}|^{2}$
$+\sum_{l=l_{0}+2i}^{l_{0}+2i+a_{0}-1}\sum_{\alpha=1}^{n_{l}}\int_{P_{l}^{(\alpha)}}|\nabla u_{l,\alpha}|^{2}+\sum_{l=l_{0}+2i+a_{0}}^{l_{0}+3i+a_{0}-1}\sum_{\alpha=1}^{n_{l}}\phi_{2}^{2}(lL)\int_{P_{l}^{(\alpha)}}|\nabla u_{l,\alpha}|^{2}$
By Proposition 5.1, we set $\epsilon=\frac{1}{n_{l}(3i+a_{0})^{2}}$ in each sum. Therefore, we obtain
$\int_{M}|\nabla\varphi_{s_{0},a_{0}}|^{2}\leq 4C\frac{1}{(iL)^{2}}2i+4C\frac{1}{(iL)^{2}}i+\frac{1}{3i+a_{0}}\leq\frac{C}{i}.$
By weak stability, we obtain
00 0 $\leq Q(\varphi_{s_{0},a_{0}},\varphi_{s_{0},a_{0}})$
$=\int_{M}|\nabla\varphi_{s_{0},a_{0}}|^{2}-\int_{M}w\varphi_{s_{0},a_{0}}^{2}$
$\leq C/i-\int_{B_{l_{0}L}(p)}w$
$\leq C/i-T.$
Letting $i$ tend to the infinity leads to a contradiction. Thus $w(x)$ has to vanish identically. That is, $A_{M}\equiv 0$ and $\operatorname{Ric}_{X}(\eta,\eta)=0$ along $M$, completing the proof.
On the one hand, the above proof also works for weakly stable CMC hypersurface. On the other hand, as mentioned before it can be adjusted to prove Proposition 1.13.
5.2. Proof of Theorem 1.12
In this subsection, we prove Theorem 1.12. We emphasize that the mean curvature used in the sequel is the trace of the second fundamental form. First, we prove the following.
@@
A complete constant mean curvature hypersurface $M^{4}$ in $(X^{5},g)$ with the bounded geometry satisfying universal subexponential end growth condition,
(5.1) $\frac{H_{M}^{2}}{4}+\inf_{X}\operatorname{Ric}_{X}\geq\delta,$
and
(5.2) $\inf_{X}k\operatorname{-triRic}+\frac{2k^{2}+k-2}{8k-6}H_{M}^{2}\geqslant 2\epsilon_{0}$
for positive constants $\delta$, $\epsilon_{0}$ and $k\in[\frac{3}{2},2]$ as well as
$\operatorname{Ind}<\infty$
must be compact.
@@
The proof follows the same strategy in [Hong24] by the first author. For completeness we include details. Indeed, on the contrary, we assume that there exists a complete noncompact hypersurface $M$ in $X^{5}$ satisfying finite index condition, universal subexponential end growth condition and the curvature conditions in the theorem.
Since $M$ has finite index, in any end $E$ of $M$ (outside a large compact subset) there holds
[…]
$\geq\left(\frac{H_{M}^{2}}{4}+\inf_{X}\operatorname{Ric}_{X}\right)\int_{E}\varphi^{2}$
$\geq\delta\int_{E}\varphi^{2}$
for any $\varphi\in C^{\infty}_{c}(E)$. It follows from [peterwangspectrum] that if $E$ is a parabolic end, it satisfies exponential volume decay, thus has finite volume. However, in a Riemannian manifold with bounded geometry, any complete noncompact CMC hypersurface (in fact, any end of this hypersurface) has infinite volume. This contradiction implies that $M$ has no parabolic end. Thus all ends of $M$ are nonparabolic. Notice that the number of nonparabolic ends may not be finite.
Again for a finite index CMC hypersurface, there exists a compact subset such that the hypersurface is strongly stable outside (see [Fischer-Colbrie-On-complete-minimal]). Take a nonparabolic end $E$ with respect to the compact subset such that $E$ is stable. Consider its universal covering $\tilde{E}$, which is still a stable CMC hypersurface with boundary in $X^{4}$ satisfying subexponential end growth condition by the assumption.
On the one hand, by Lemma 5.1 in [Hong24], $\tilde{E}$ is nonparabolic. According to Corollary 3.6 and Lemma 3.7 in [Hong24], $\tilde{E}$ satisfies exponential volume growth:
$|\tilde{E}\cap(B_{(l+1)L}\setminus B_{lL})|\geq C\cdot e^{Cl}$
for a constant $C=C(\delta)$ depending on $\delta.$ On the other hand, by Lemma 4.8 each component of $\tilde{E}\cap(B_{(l+1)L}\setminus B_{lL})$ has uniform volume upper bound. It follows from the subexponential end growth condition that the volume of $\tilde{E}\cap(B_{(l+1)L}\setminus B_{lL})$ satisfies
$|\tilde{E}\cap(B_{(l+1)L}\setminus B_{lL})|=o(e^{cl})$
for any $c>0.$ This contradicts with the exponential volume growth. ∎
Proof of Theorem 1.12.
We only need to check the curvature conditions in Theorem 5.2. The first one leads to that $H_{M}>4$. The second one turns to finding the minimum of $H_{M}$ satisfying the followng inequality in terms of $k\in[\frac{3}{2},2]$:
$H_{M}^{2}>\frac{(8k-6)(4k+5)}{2k^{2}+k-2}.$
When $k=2,$ the right hand side achieves the minimum $65/4$
@@ Appendix A Example
We provide a closed $5$-dimensional Riemannian manifold $(X^{5},g)$ with bounded geometry, satisfying
$\operatorname{Ric}_{X}\geq 0,\ \,\operatorname{biRic}_{X}\geq 0,\ \,\text{and}\ \ 2\operatorname{-triRic}_{X}\geq\delta>0,$
@@
Note that in general,
$\operatorname{Ric}_{g_{s}}(v,v)\leq\operatorname{Ric}_{g_{s}}(e_{2},e_{2})=\operatorname{Ric}_{g_{s}}(f_{1},f_{1})=\operatorname{Ric}_{g_{s}}(f_{2},f_{2})\leq\operatorname{Ric}_{g_{s}}(\overline{e}_{1},\overline{e}_{1}).$
@@
$2\operatorname{-triRic}(x;y,z)=2\operatorname{Ric}(x,x)+\operatorname{biRic}(y,z)-\sec(x,y)-\sec(x,z).$
Therefore, if $02\operatorname{Ric}(x,x)-\sec(x,y)-\sec(x,z)>$ 0, we have $02\operatorname{-triRic}(x;y,z)>$ 0 automatically.
Direct calculation yields when $s^{2}\in(16/9,11/3]$
$\begin{split}2\operatorname{-triRic}(v;\overline{e}_{1},e_{2})=2\operatorname{-triRic}(v;\overline{e}_{1},f_{2})&\geq 11+\frac{3}{2}s^{2}+6\left(1-\frac{3}{4}s^{2}\right)>0,\\ 2\operatorname{-triRic}(v;\overline{e}_{1},f_{1})>0,\quad&2\operatorname{-triRic}(v;e_{2},f_{1})>0,\\ 2\operatorname{-triRic}(v;e_{2},f_{2})\geq 18-\frac{15}{4}s^{2}>0,\quad&2\operatorname{-triRic}(v;f_{1},f_{2})>0.\end{split}$
$\begin{split}2\operatorname{-triRic}(\overline{e}_{1};v,e_{2})=2\operatorname{-triRic}(\overline{e}_{1};v,f_{2})&\geq 17-3s^{2}>0,\\ 2\operatorname{-triRic}(\overline{e}_{1};v,f_{1})>0,\quad&2\operatorname{-triRic}(\overline{e}_{1};e_{2},f_{1})>0,\\ 2\operatorname{-triRic}(\overline{e}_{1};e_{2},f_{2})\geq 20-\frac{15}{4}s^{2}>0,\quad&2\operatorname{-triRic}(\overline{e}_{1};f_{1},f_{2})>0,\end{split}$
$\begin{split}2\operatorname{-triRic}(e_{2};v,\overline{e}_{1})=17-3s^{2}>0,\quad&2\operatorname{-triRic}(e_{2};v,f_{1})\geq 18-\frac{15}{4}s^{2}>0,\\ 2\operatorname{-triRic}(e_{2};v,f_{2})>0,\quad&2\operatorname{-triRic}(e_{2};\overline{e}_{1},f_{1})\geq 18-\frac{15}{4}s^{2}>0,\\ 2\operatorname{-triRic}(e_{2};\overline{e}_{1},f_{2})>0,\quad&2\operatorname{-triRic}(e_{2};f_{1},f_{2})>0.\end{split}$
$\begin{split}2\operatorname{-triRic}(f_{1};v,\overline{e}_{1})=17-3s^{2}>0,\quad&2\operatorname{-triRic}(f_{1};v,e_{2})>0,\\ 2\operatorname{-triRic}(f_{1};v,f_{2})\geq 21-6s^{2},\quad&2\operatorname{-triRic}(f_{1};\overline{e}_{1},e_{2})\geq 18-\frac{15}{4}s^{2}>0,\\ 2\operatorname{-triRic}(f_{1};\overline{e}_{1},f_{2})>0,\quad&2\operatorname{-triRic}(f_{1};e_{2},f_{2})\geq 20-\frac{15}{4}s^{2}>0.\end{split}$
To obtain $02\operatorname{-triRic}(f_{1};v,f_{2})>$ 0, we require $s^{2}<7/2<11/3$.
$\begin{split}2\operatorname{-triRic}(f_{2};v,\overline{e}_{1})=17-3s^{2}>0,\quad&2\operatorname{-triRic}(f_{2};v,e_{2})>0,\\ 2\operatorname{-triRic}(f_{2};v,f_{1})\geq 18-\frac{15}{4}s^{2}>0,\quad&2\operatorname{-triRic}(f_{2};\overline{e}_{1},e_{2})>0,\\ 2\operatorname{-triRic}(f_{2};\overline{e}_{1},f_{1})\geq 20-\frac{15}{4}s^{2}>0,\quad&2\operatorname{-triRic}(f_{2};e_{2},f_{1})>0.\end{split}$
In conclusion, when $216/9<s^{2}<7/$ 2, $(\mathbb{S}^{5},g_{s})$ has positive $\operatorname{Ric}$, $\operatorname{biRic}$ and $2\operatorname{-triRic}$ curvatures with some negative sectional curvature.
However, when $210/3<s^{2}<7/$ 2,
@@
which implies that $\operatorname{Ric}_{3}<0$ somewhere. By the main theorem, such ambient manifold does not admit any two-sided complete noncompact minimal hypersurfaces.
Moreover, note that when $s^{2}\in(\frac{11}{3},4)$, $\operatorname{Ric}_{g_{s}}\geq 3\left(2-\frac{s^{2}}{2}\right)>0$, while $\operatorname{biRic}_{g_{s}}(e_{1},e_{2})=11-3s^{2}<0$ and $\inf\sec_{g_{s}}=4-\frac{9}{4}s^{2}<0$. It is interesting to understand whether there exists a two-sided complete noncompact stable minimal hypersurface in such manifold $(\mathbb{S}^{5},g_{s})$.

### v3

@@ 5. Proof of main results
We are going to prove a more general result than ones stated in the introduction. For clarity, we state our result and divide proofs into different cases.
5.1. Case 1: $n=4$
Theorem 5.1.
Let $(X^{5},g)$ be a complete manifold and let $M^{4}\hookrightarrow(X^{5},g)$ be a complete noncompact stable CMC immersion with mean curvature $H$.
$(1)$ If $\operatorname{Ric}^{X}+\frac{1}{4}H^{2}\geq 0$, $\operatorname{biRic}_{X}+\frac{1}{4}H^{2}\geq 0$, $X$ has bounded geometry, and $\lambda_{k}+\frac{2k^{2}+k-2}{8k-6}H^{2}\geq 2\varepsilon_{0}$ for a positive constant $\varepsilon_{0}$ and some $k\in[1,2]$. Then $M$ is parabolic with at most two ends. In particular, $M$ is umbillic and $\operatorname{Ric}^{X}(\nu,\nu)+\frac{1}{4}H^{2}=0$ along $M$.
$(2)$ If $\operatorname{Ric}^{X}_{3}+\frac{3}{28}H^{2}\geq 0$, $X$ has weakly bounded geometry, and $\lambda_{k}+\frac{2k^{2}+k-2}{8k-6}H^{2}\geq 2\varepsilon_{0}$ for a positive constant $\varepsilon_{0}$ and some $k\in[1,2]$. Then $M$ is parabolic with at most two ends. In particular, $M$ is umbillic and $\operatorname{Ric}^{X}(\nu,\nu)+\frac{1}{4}H^{2}=0$ along $M$.
Remark 5.2.
Note that $\operatorname{Ric}^{X}_{3}+\frac{3}{28}H^{2}\geq 0$ implies $\operatorname{Ric}^{X}+\frac{1}{4}H^{2}\geq 0$ and $\operatorname{biRic}_{X}+\frac{1}{4}H^{2}\geq 0$.
Proof.
First, observe that the universal cover of $M$ is stable, so without loss of generality, we may assume that $M$ is simply connected. The theorem then follows by constructing suitable compactly supported test functions. To achieve this, we decompose $M$ into several geometric building blocks, following the approach in [chodosh-li-stryker]. Let $M^{n}$ be a complete, simply connected Riemannian $n$-manifold.
(1): It follows from Theorem 3.1 that $M$ has at most two ends when $\operatorname{biRic}_{X}+\frac{1}{4}H^{2}\geq 0$. Furthermore, if there exist more than two ends, $M$ is a product manifold with two ends. In the sequel, we only need to address the one end case. We denote $\{E_{l}\}$ the unique end of $M$ adapted to $p\in M$ with length scale $L$.
We now demonstrate the decomposition of $M$. For fixed $l_{0}\geqslant 1$, we set
@@ 5.1. Case 1: $n=4$
By the uniqueness of the end, all components $P_{l}^{(\alpha)}$ are bounded. It also should be noted that the number $n_{l}$ of components $P_{l}^{(\alpha)}$ may be unbounded as $l$ tends to infinity. For fixed $i\geq 1$, we have the decomposition (see Figure 1)
We now construct the desired cutoff function. For each $l$, let $\rho_{l}$ be a smooth function on $\overline{M}_{l}$ such that $|\nabla\rho_{l}|\leqslant 2$,
For instance, we can take $\rho_{l}$ to be a smoothing of the distance function from $p$. Let $\phi:[l_{0}L,(l_{0}+i)L]\to[0,1]$ be the linear function defined by
$\phi(t)=1+\frac{1}{i}(l_{0}-\frac{t}{L}).$
Define a compactly supported Lipschitz function on $M$ as follows:
$\varphi(x)=\begin{cases}1\ \ \ \ \ \ \ \ \ \ &x\in\overline{B}_{l_{0}L}(p),\\ \phi(\rho_{l}(x))&x\in\overline{M_{l}},\ \ \l_{0}\leq l<l_{0}+i,\\ \phi(lL)&x\in\overline{P}_{l}^{(\alpha)},\ \ l_{0}\leq l<l_{0}+i,\\ 0&\text{otherwise}.\end{cases}$
For any positive superharmonic function $\omega$ on $M$, we define $\tau=\log\omega$ which satisfies $\Delta\tau\leq-|\nabla\tau|^{2}.$ Integration by parts and Cauchy-Schwarz inequality yield
$\int_{M}|\nabla\tau|^{2}\varphi^{2}$ $\leq-\int_{M}\varphi^{2}\Delta\tau$
$=\int_{M}2\varphi\nabla\varphi\cdot\nabla\tau$
$\leq\delta\int_{M}|\nabla\varphi|^{2}+\frac{1}{\delta}\int_{M}|\nabla\tau|^{2}\varphi^{2}.$
Choosing $\delta>1$, we have
$\int_{M}|\nabla\tau|^{2}\varphi^{2}$ $\leq C\int_{M}|\nabla\varphi|^{2}$
$\leq C\sum_{l=l_{0}}^{l_{0}+i-1}\int_{\bar{M}_{l}}\frac{1}{(iL)^{2}}|\nabla\rho_{l}|^{2}$
$\leq C\frac{1}{iL^{2}}\to 0,\quad i\to 0.$
Here we have used the volume estimate in Theorem 4.9 in the last inequality. Thus, both $\tau$ and $\omega$ are constant functions, and $M$ is parabolic.
Then, by the stability of $M$, there is a positive superharmonic function $\omega$ on $M$ satisfying
$\Delta\omega=-|A|^{2}\omega-\operatorname{Ric}^{X}(\nu,\nu)\omega\leq-\frac{1}{4}H^{2}\omega-\operatorname{Ric}^{X}(\nu,\nu)\omega\leq 0.$
By previous discussion, we conclude that $M$ is umbillic and has $\operatorname{Ric}^{X}(\nu,\nu)+\frac{1}{4}H^{2}=0$ along $M.$
@@
(2): By Remark 5.2 and Theorem 3.1, $M$ has at most two ends. Then the proof is exactly same as the one in (1). Notice that we can apply Lemma 4.7 and Theorem 4.9 under the assumptions of (2).
5.2. Case 2: $n=5$
@@
Let $(X^{6},g)$ be a complete manifold and let $M^{5}\hookrightarrow(X^{6},g)$ be a complete noncompact stable CMC immersion with mean curvature $H$.
$(1)$ If $\operatorname{Ric}^{X}+\frac{1}{5}H^{2}\geq 0$, $\operatorname{biRic}_{X}\geq 0$, $X$ has bounded geometry, and $\lambda_{(0.99,1.001)}+\delta(0.99,5,1.001)H^{2}\geq 2\varepsilon_{0}$ for a positive constant $\varepsilon_{0}$ where $\delta$ is taken from Proposition 4.6. Then $M$ is parabolic. In particular, $M$ is umbillic and $\operatorname{Ric}^{X}(\nu,\nu)+\frac{1}{5}H^{2}=0$ along $M$.
$(2)$ If $\operatorname{Ric}^{X}_{4}\geq 0$, $H=0$, $X$ has weakly bounded geometry, and $\lambda_{(0.99,1.001)}\geq 2\varepsilon_{0}$ for a positive constant $\varepsilon_{0}$. Then $M$ is parabolic. In particular, $M$ is totally geodesic and $\operatorname{Ric}^{X}(\nu,\nu)=0$ along $M$.
@@
(1) and (2): Under current assumptions, there is no priori upper bound on the number of ends. Fortunately, by Theorem 3.3, we know that the number of nonparabolic end is at most one if either $\operatorname{biRic}_{X}\geq 0$ and $X$ has bounded geometry or $\operatorname{Ric}^{X}_{4}\geq 0$ and $M$ is minimal.
Then the rest of proof is basically same as that of Theorem 5.1 (1), except that the component $\overline{P}_{k}^{(\alpha)}$ now is not necessarily bounded. On each unbounded parabolic component $\overline{P}_{k}^{(\alpha)}$, the modified test function is given by the following well-known fact: for any $l$ and $\alpha\in[1,n_{l}]$ and for any small $\epsilon>0$, there exists a positive compactly supported Lipschitz function $u_{l,\alpha}$ on $P_{l}^{(\alpha)}$ such that
(5.1) $u_{l,\alpha}|_{\partial P_{l}^{(\alpha)}}=1\ \text{and}\ \ \int_{P_{l}^{(\alpha)}}|\nabla u_{l,\alpha}|^{2}<\epsilon.$
This fact follows from the definition of parabolic ends (see [peterwangspectrum] or [chodosh-li-stryker, Proposition 3.4] for more details). Thus on $\overline{P}_{l}^{(\alpha)}$, we now define $\varphi:=\phi(lL)u_{l,\alpha}$. Besides, on bounded components, we set $\varphi:=\phi(lL)$ as we previously did.
We can choose $\epsilon$ sufficiently small in (5.1) such that the energy of all parabolic ends on a compact subset in $M$ is sufficiently small. Moreover, under current assumptions, we can apply Lemma 4.7 and Theorem 4.9.
@@ Appendix A Example
We provide a (in fact, a family of) closed $5$-dimensional Riemannian manifold $(X^{5},g)$ with bounded geometry, satisfying
$\operatorname{Ric}_{X}\geq 0,\ \,\operatorname{biRic}_{X}\geq 0,\ \,\text{and}\ \ 1\operatorname{-triRic}_{X}\geq\delta>0,$
@@
$1\operatorname{-triRic}(x;y,z)=\operatorname{Ric}(x,x)+\operatorname{biRic}(y,z)-\sec(x,y)-\sec(x,z),$
and note that $1\operatorname{-triRic}(x;y,z)$ is symmetric in $(x,y,z)$.
Direct calculation yields when $s^{2}\in(16/9,11/3]$,
$1\operatorname{-triRic}(v;\overline{e}_{1},e_{2})$ $=1\operatorname{-triRic}(v;\overline{e}_{1},f_{2})=11-\frac{3}{2}s^{2}>0,$
$1\operatorname{-triRic}(v;\overline{e}_{1},f_{1})$ $=8+\frac{3}{4}s^{2}>0,$
$1\operatorname{-triRic}(v;e_{2},f_{1})$ $=1\operatorname{-triRic}(v;f_{1},f_{2})=11-\frac{3}{2}s^{2}+(1-\frac{3}{4}s^{2})\sin^{2}\alpha>0,$
$1\operatorname{-triRic}(v;e_{2},f_{2})$ $=8+\frac{3}{4}s^{2}+(4-3s^{2})\sin^{2}\alpha>0;$
$1\operatorname{-triRic}(\overline{e}_{1};e_{2},f_{1})$ $=1\operatorname{-triRic}(\overline{e}_{1};f_{1},f_{2})=12-\frac{9}{4}s^{2}+(-1+\frac{3}{4}s^{2})\sin^{2}\alpha>0,$
$1\operatorname{-triRic}(\overline{e}_{1};e_{2},f_{2})$ $=12-\frac{9}{4}s^{2}+(3s^{2}-4)\sin^{2}\alpha>0;$
$1\operatorname{-triRic}(e_{2};f_{1},f_{2})$ $=12-\frac{9}{4}s^{2}>0.$
In conclusion, when $316/9<s^{2}<11/$ 3, $(\mathbb{S}^{5},g_{s})$ has positive $\operatorname{Ric}$, $\operatorname{biRic}$ and $1\operatorname{-triRic}$ curvatures with some negative sectional curvature.
Moreover, when $10/3<s^{2}$,
@@
which implies that $\operatorname{Ric}_{3}<0$ somewhere. Thus, $(\mathbb{S}^{5},g_{s})$ with $s\in(10/3,11/3)$ is a family of desired ambient manifolds. By Theorem 1.9, such ambient manifolds do not admit any two-sided complete noncompact minimal hypersurfaces.
To end the appendix, we would like to mention that when $s^{2}\in(\frac{11}{3},4)$, $\operatorname{Ric}_{g_{s}}\geq 3\left(2-\frac{s^{2}}{2}\right)>0$, while $\operatorname{biRic}_{g_{s}}(e_{1},e_{2})=11-3s^{2}<0$ and $\inf\sec_{g_{s}}=4-\frac{9}{4}s^{2}<0$. This means that positive Ric curvature ($C_{1}$ curvature) does not imply positive $\operatorname{biRic}$ ($C_{2}$ curvature) in general. It is also interesting to understand whether there exists a two-sided complete noncompact stable minimal hypersurface in such manifolds $(\mathbb{S}^{5},g_{s})$.